A 175.0 g sample of ice at 0.0 ̊C ends up as liquid water at 45.0 ̊C. How much energy is involved in this change?
1
Expert's answer
2020-12-14T14:34:43-0500


Two processes occur here that require energy. (i) Phase change(melting), and (ii) temperature change;

I. Q during Phase change


Q1=mLf


"=\dfrac{175.0}{1000}Kgx334J/Kg"

= 58.45J


II. Q for temperature change

Q2=mC"\Delta"T

"=175.0gx\dfrac{4.18J}{g^oC} x 45.0^oC"

= 32,917.5J


Add up for total energy

QT=Q1 + Q2

= 58.45J + 32,917.5J

= 32,975.98J

= 32.98kJ